Prev Next

SW&IT Meeting Minutes 2022-01-03

Attendees: Brandon, David, Glenn, Matthieu, Paul, Stephen, Xianyu, Yang

Outlook

Labor Reporting

Report auto-generated every Monday at 6:30am, Do not edit this table, Update your entries in smartsheets.

News from management

  • director search, update from AF

Previous Action Items

Owner Action Due Date Completion Date

Open topics

Team reports

MB

  • Last week
    • vacation
  • This week
    • resume dms lit2
    • aob

GE

  • Two weeks ago
    • Short week due to holidays.
    • Released the test version of the new RTD to SciOps on observer@obs5. A test plan can be found in my Google Docs.
  • This week
    • Begin working on the changes to the RTD needed for subwindowing.
    • Merge the AO fix (IT 8530) and deploy to luci.luci.
    • Merge the new RTD, and deploy to the observer computers.
    • Check for use of Log4J and deploy Log4J update.
    • Examine the other LUCI GUIs to check for GUI modifications from threads other than the Event Dispatch Thread.

DE

  • Louise desktop situation
  • Synology drives
    • Followup on SSD. Argh.
  • Looks like the calendar reminder for the archive directory creation headed off our annual panic. Yay!
  • Ketiv meeting
  • Safety training
  • Various backlog stuff will come up I'm sure

BM

  • Two weeks ago, small fixes:
    • Python AOApps show the right name in log files
    • CenterPupils bestEffort=true if WFS is LBTI
    • Figured out more about wrong filotto issue (why they don't show up in .bins)--comment on Github.
  • This week:
    • Ramp back up on PB upgrade – where are we, what specific questions do we have for GB/XZ.
    • More "diagnostic" issues: masterdiagnostic field width, vague anemometer error logging, etc.
    • "Annoying" issues: manual safeskip alert button, clear gopt plots when new preset is loaded, etc.
    • LBTI SafeSkip problem needs to be reproduced with LBTI people. Manual SafeSkip alert button should alleviate annoyance.
    • Vacation next Tuesday-Friday (Jan 11-14)

SH

  • Some calls but nothing major (network, overhead displays, PLC stuff).

next:

  • VMware (maybe)
  • Overhead displays
  • Some edits to the info page
  • New sciops website (check in)

people should be aware of:
  • Weather on the mountain
  • network related to weather on the mountain
  • telescope related to weather on the mountain

PSG

  • Last week
    • Vacation/New Year's
  • This week
    • Finish implementing changes per comments on pull request.
    • Update pull request.
    • Create schem file for "Jenkins" for Argos Proxy.
Topic revision: r9 - 04 Jan 2022, PaulGrenz
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback